OpenStudy (anonymous):

Consider the functions f(x) = 2(x − 9) and g(x) = 4x. What is the value of (f + g)(5)? −8 28 20 12

OpenStudy (anonymous):

(f+g)(5)=2(x-9)+4x=2x-18+4x=6x-18=6(5)-18=30-18=12